What are the values of r in the equation r2-7r-8=0
Answer by ewatrrr(24785)


Hi,
Solving for r
r^2-7r-8=0
Factoring
(r - 8)(r + 1)=0 Note:SUM of the inner product(-8r) and the outer product(1*r) = -7r
(r - 8)= 0
r = 8
(r + 1)=0
r = -1